# xGen-MM-Vid (BLIP-3-Video): You Only Need 32 Tokens to Represent a Video Even in VLMs

### v1

@@ Abstract
We present xGen-MM-Vid (BLIP-3-Video): a multimodal language model for videos, particularly designed to efficiently capture temporal information over multiple frames. BLIP-3-Video takes advantage of the ‘temporal encoder’ in addition to the conventional visual tokenizer, which maps a sequence of tokens over multiple frames into a compact set of visual tokens. This enables BLIP-3-Video to use much fewer visual tokens than its competing models (e.g., 32 vs. 4608 tokens). We explore different types of temporal encoders, including learnable spatio-temporal pooling as well as sequential models like Token Turing Machines. We experimentally confirm that BLIP-3-Video obtains video question-answering accuracies comparable to much larger state-of-the-art models (e.g., 34B), while being much smaller (i.e., 4B) and more efficient by using fewer visual tokens. The project website is at https://www.salesforceairesearch.com/opensource/xGen-MM-Vid/index.html
@@ 1 Introduction
Further, in addition to such VLMs trained with images, VLMs for videos are becoming increasingly popular. The key component in a VLM for videos is the temporal abstraction of tokens over multiple frames. Models like Video-ChatGPT (Maaz et al., 2024) and PLLaVA (Xu et al., 2024a) rely on a simple spatial/temporal pooling on top of image frame-level tokens to represent the entire video. Some models rely on a separate video encoder to capture temporal information in videos (Lin et al., 2023). Similarly, some models use additional convolutional layers (or Transformer layers) over frames to reduce their representation size (e.g., Video-LLaMA (Zhang et al., 2023), Kangaroo (Liu et al., 2024)). Approaches that simply collect all the visual tokens from all the frames (e.g., MiniGPT4-video (Ataallah et al., 2024), LLaVA-NeXT (Li et al., 2024b), Tarsier (Wang et al., 2024a) and LLaVA-OneVision (Wang et al., 2024a)) also have been very popular recently, as they allow capturing all the details from the frame-level tokens.
Figure 1: SOTA video VLM model comparison: (Left) Number of visual tokens vs. video-QA accuracy. (Right) Model size vs. video-QA accuracy.
@@
In this paper, we introduce xGen-MM-Vid (BLIP-3-Video), which is an efficient compact vision-language model with an explicit temporal encoder, designed particularly for videos. BLIP-3-Video particularly focuses on incorporating a learnable ‘temporal encoder’ within it. We explore different types of temporal encoder, and demonstrate that the model can abstract each video into much fewer visual tokens (e.g., 16) while being successful in open-ended question-answering and captioning tasks. We include a space-time attentional pooling as well as a sequential model as our temporal encoder, relying on token operations to iteratively abstract a series of frame-level tokens into a learnable memory.
There has been prior work investigating the role of pooling (Jin et al., 2024), convolutions, and cross attention layers (Zhang et al., 2023; Liu et al., 2024; Li et al., 2024c), but study on full space-time attentional pooling or sequential model to this extent has been limited in the past. Our objective in this paper is to provide a fundamental alternative to more brute-force way of collecting all the visual tokens which have been increasing popular recently. We experimentally confirm that $16\sim 32$ video tokens abstracted by the temporal encoder is often sufficient to represent the entire video for question-answering (Figure 1).
@@ 2.1 Model architecture
Figure 2: An illustration of the BLIP-3-Video model architecture. It has the explicit temporal encoder inserted to BLIP-3.
We build the BLIP-3-Video model based on the image-based vision-language model (VLM), BLIP-3 (Xue et al., 2024).
The model architecture is composed of the following four components: (1) the vision encoder (ViT) taking each frame input, (2) the frame-level tokenizer to reduce the number of tokens, (3) the temporal encoder to build video-level token representations, and (4) the autoregressive LLM generating output text captions based on such video tokens and text prompt tokens. Figure 2 shows an overview.
@@
We use Phi-3 (Abdin et al., 2024) as our LLM backbone taking such video tokens in addition to the text prompt tokens. This enables the model to take text+video as an input and generate text sentences as an output.
@@ 2.2 Temporal encoders
A temporal encoder is a function of tokens, taking $N\cdot T$ tokens as an input and returning $M$ tokens as an output: $x_{1,\ldots,M}=f(v_{(1,1),\ldots,(N,T)})$.
We explore different types of encoders as part of our model. The simplest form of the temporal encoder will be temporal pooling, e.g., summating per-frame tokens over time: $x_{1,\ldots,M}=\left\{\sum_{t}(v_{(i,t)})\right\}_{i=1}^{M}$ where $M$ is always restricted to be identical to $N$, which was also used in (Maaz et al., 2024). Another possible implementation would be the use of a temporal Transformer, modeling the entire token sequence and selecting the last $m$ tokens similar to Mirasol3B (Piergiovanni et al., 2024):
@@
In addition to the straightforward temporal encoders mentioned above, we explore two important temporal encoders considering space-time nature of tokens: spatio-temporal attentional pooling and sequential models (Figure 3).
Spatio-temporal attentional pooling:
Attentional pooling allows learnable ‘soft selection’ of multiple tokens given a larger set of tokens. Such attentional pooling have been previously developed for Transformers (e.g., Perceiver (Jaegle et al., 2022) and TokenLearner (Ryoo et al., 2021)), and also used in earlier foundation models (e.g., CoCa (Yu et al., 2022)) for images.
In our model, we use TokenLearner (Ryoo et al., 2021), making it explicitly serve as our space-time aware temporal encoder. Unlike previous per-image-frame usage of poolings where spatial pooling and temporal pooling are applied separately (e.g., Video-ChatGPT), our temporal encoder directly takes all $N\cdot T$ tokens and ‘learns’ to soft-select $M$ informative tokens spatio-temporally. Here, $N$ tokens could be viewed as spatial representations of a frame and we have $T$ of them, suggesting it is a spatio-temporal representation selection.
@@
Sequential Model:
We also deploy Token Turing Machines (TTM) (Ryoo et al., 2023) as a temporal encoder, which is a sequential model capable of taking any number of frames to generate a video-level token representation (e.g., $M=32$ regardless the number of frames). Our use of TTM is similar to its usage in Mirasol3B (Piergiovanni et al., 2024), except that our model uses TTM directly to encode a sequence of image tokens while Mirasol3B uses TTM to encode a sequence of sets of video tokens.
Here, we also further extend TTM by adding time-stamped positional encodings to embded the frame index of each token in the latent space. This enables the tokens in the ‘memory’ of TTM to preserve the temporal ordering information, which is crucial when representing complicated or long video scenes. In addition, we use TTM temporal encoder in a ‘grouped’ fashion, maintaining a separate memory of size $G=4$ for each of $N=128$ tokens over time. The memory is maintained to have the size is $N\cdot G$, and the final output from the sequence model is attentionally pooled from the final memory to give $M$ tokens.
@@ 2.3 Training recipe
BLIP-3-Video follows a three-stage curriculum learning: (1) image caption pretraining, (2) video caption pretraining, and (3) video instruction tuning. In all its training we freeze the vision encoder, only training the model parameters after the vision encoder. First, we directly use the pretrained weights from BLIP-3 (Xue et al., 2024). BLIP-3 is for images and it does not contain weights for the temporal encoder, so we randomly initialize those weights.
As its 2nd stage, the model is then trained on LLaVA-Hound-DPO’s video caption data (Zhang et al., 2024b), featuring over 900k video captions. Instead of directly using the text captions provided in LLaVA-Hound-DPO, we used GPT-4 to rephrase such text captions so that they become more GPT-style captions.
Finally, we tuned the model using a mix of video question-answering datasets, including VideoChatGPT’s 99k-sample video instruction tuning data (Maaz et al., 2024), along with the training splits of the MSVD-QA (Xu et al., 2017), MSRVTT-QA (Xu et al., 2017), ActivityNet-QA (Yu et al., 2019), TGIF-QA (Jang et al., 2017), and NExT-QA (Xiao et al., 2021) datasets, which contain 30k, 149k, 32k, 71k, and 34k samples, respectively. For TGIF-QA, we only used the training data associated with the Repeating Action and State Transition tasks. In our video instruction tuning recipe, we employ both open-ended and multiple-choice video QA formats for TGIF-QA and NExT-QA. For the open-ended video QA training data sourced from the MSVD-QA, MSRVTT-QA, TGIF-QA, and NExT-QA training sets, we used GPT-3.5 to rephrase the original single-word or single-phrase answer into a natural language sentence, providing the question in the LLM prompt context. For open-ended TGIF-QA and NExT-QA, we also double the sample size by using both the original short-phrase answers and the rephrased sentence-based answers. In addition, we added a filtered version of the Mira caption dataset (Ju et al., 2024) for our video instruction tuning. That is, we are using both video question-answering and video captioning for our final training. We excluded captions for Mira videos longer than one minute, totaling 935k video caption samples.
@@ 3 Experiments and Results
3.1 Model implementation details
We share the model details with BLIP-3 (4B), except that BLIP-3-Video has the new temporal encoder component in its architecture. This model takes the video with the input resolution of 384$\times$384, using SigLIP encoder to map it to 729 tokens per frame with the channel size 1152. Perceiver-Resampler is implemented with multiple cross-attention layers with the same channel dim, which is then given to the temporal encoder.
TokenLearner serving as the spatio-temporal attentional pooling was implemented using a MLP as the attention function. The size of its inner dim was the number of target tokens * 2. The grouped TTM serving as the sequential model temporal encoder was implemented using 4 Transformer layers (with the channel dim of 1152) as the processor module while using TokenLearners for read/write modules. Memory size was set to $N*4=512$ tokens total.
The resulting $16\sim 128$ tokens are mapped to the text embedding with the channel dimension of 3072, before given to the LLM (Phi-3).
3.2 Public benchmarks
We conducted experiments measuring video question-answering accuracies on multiple public datasets. This includes open-ended answer generation tasks like MSVD-QA, as well as multiple choice questions like NExT-QA. We follow their standard settings in all cases.
With the temporal encoder, BLIP-3-Video was able to retain the performance with much fewer tokens, which we discuss more in the following subsection. Our results suggest that not too many visual tokens are really necessary to be successful on these open-ended question answering benchmarks, as long as we have a carefully designed temporal encoder.
In addition, we evaluated BLIP-3-Video’s ability to solve multiple choice questions (MCQ). Table 2 shows the results on NExT-QA. Due to the nature of its questions requiring understanding of multiple frames, many prior models use quie a bit of tokens. For instance, GPT-4 uses a minimum of 255 tokens per frame. It is interesting that BLIP-3-Video achieves comparable accuracy while representing the entire video with only 32 (or 128) tokens.
@@ 3.2 Public benchmarks
VideoChat (Li et al., 2023b) 7B 32 56.3 / 2.8 45.0 / 2.5 - / 2.2 34.4 / 2.3
Video-LLaMA (Zhang et al., 2023) 7B 32 51.6 / 2.5 29.6 / 1.8 12.4 / 1.1 - / -
Video-ChatGPT (Maaz et al., 2024) 7B 264+ 64.9 / 3.3 49.3 / 2.8 34.2 / 2.8 51.4 / 3.0
Chat-UniVi (Jin et al., 2024) 7B 112 69.3 / 3.7 55.0 / 3.1 46.1 / 3.3 69.0 / 3.8
LLaMA-VID (Li et al., 2024c) 7B 32 69.7 / 3.7 57.7 / 3.2 47.4 / 3.3 -
LLaMA-VID (Li et al., 2024c) 13B 32 70.0 / 3.7 58.9 / 3.3 47.5 / 3.3 -
Video-LLaVA (Lin et al., 2023) 7B 2048 71.8 / 3.9 59.2 / 3.5 45.3 / 3.3 70.0 / 4.0
MiniGPT4-Video (Ataallah et al., 2024) 7B 2880+ 73.9 / 4.1 59.7 / 3.3 46.3 / 3.4 72.2 / 4.1
PLLaVA (Xu et al., 2024a) 7B 576+ 76.6 / 4.1 62.0 / 3.5 56.3 / 3.5 77.5 / 4.1
SlowFast-LLaVA Xu et al. (2024b) 7B 3680 79.1 / 4.1 65.8 / 3.6 56.3 / 3.4 78.7 / 4.2
LLaVA-Hound-DPO Zhang et al. (2024b) 7B 2048 80.7 / 4.1 70.2 / 3.7 - / - 61.4 / 3.5
LLaVA-OneVision (Wang et al., 2024a) 7B 1568 72.9 / 3.9 57.8 / 3.4 55.3 / 3.6 41.1 / 3.1
Tarsier (Wang et al., 2024a) 7B 4608+ 77.0 / 4.1 62.0 / 3.5 59.5 / 3.6 79.2 / 4.2
Tarsier (Wang et al., 2024a) 7B 4608 74.4 / 4.0 59.1 / 3.4 54.3 / 3.5 - / -
PLLaVA (Xu et al., 2024a) 34B 576+ 79.9 / 4.2 68.7 / 3.8 60.9 / 3.7 80.6 / 4.3
LLaVA-NeXT-Video (Li et al., 2024b) 32B 1152 73.6 / 4.0 56.8 / 3.4 58.4 / 3.6 73.5 / 4.1
Tarsier (Wang et al., 2024a) 34B 4608+ 80.3 / 4.2 66.4 / 3.7 61.6 / 3.7 82.5 / 4.4
Tarsier (Wang et al., 2024a) 34B 4608 79.3 / 4.1 62.2 / 3.5 61.5 / 3.7 - / -
@@
Method Size #tokens NExT-QA
LangRepo (Kahatapitiya et al., 2024) 7B 3136+ 54.6
LangRepo (Kahatapitiya et al., 2024) 12B 3136+ 60.9
Tarsier (Wang et al., 2024a) 7B 4608+ 71.6
LLoVi (Zhang et al., 2024a) 157B 1000s 67.7
IG-VLM (Kim et al., 2024) 34B 1536+ 70.9
VideoAgent (Wang et al., 2024b) GPT-4 2091+ 71.3
VideoTree (Wang et al., 2024c) GPT-4 3978+ 73.5
Tarsier (Wang et al., 2024a) 34B 4608+ 79.2

### v2

@@ Abstract
We present xGen-MM-Vid (BLIP-3-Video): a multimodal language model for videos, particularly designed to efficiently capture temporal information over multiple frames. BLIP-3-Video takes advantage of the ‘temporal encoder’ in addition to the conventional visual tokenizer, which maps a sequence of tokens over multiple frames into a compact set of visual tokens. This enables BLIP-3-Video to use much fewer visual tokens than its competing models (e.g., 32 vs. 4608 tokens). We explore different types of temporal encoders, including learnable spatio-temporal pooling as well as sequential models like Token Turing Machines. We experimentally confirm that BLIP-3-Video obtains video question-answering accuracies comparable to much larger state-of-the-art models (e.g., 34B), while being much smaller (i.e., 4B) and more efficient by using fewer visual tokens.
@@ 1 Introduction
Figure 1: SOTA video VLM model comparison: Number of visual tokens vs. video-QA accuracy.
In addition to VLMs trained with images, VLMs for videos are becoming increasingly popular. The key component in a VLM for videos is the temporal abstraction of tokens over multiple frames. Models like Video-ChatGPT [31] and PLLaVA [49] rely on a simple spatial/temporal pooling on top of image frame-level tokens to represent the entire video. Some models rely on a separate video encoder to capture temporal information in videos [28]. Similarly, some models use additional convolutional layers (or Transformer layers) over frames to reduce their representation size (e.g., Video-LLaMA [56], Kangaroo [30]). Approaches that simply collect all the visual tokens from all the frames (e.g., MiniGPT4-video [3], LLaVA-NeXT [23], Tarsier [43] and LLaVA-OneVision [43]) also have been very popular recently, as they allow capturing all the details from the frame-level tokens.
@@
Figure 2: BLIP-3-Video model architecture. It has an explicit temporal encoder inserted to BLIP-3.
In this paper, we introduce BLIP-3-Video, which is an efficient compact vision-language model with an explicit temporal encoder, designed particularly for videos. BLIP-3-Video particularly focuses on incorporating a learnable ‘temporal encoder’ within it. We explore different types of temporal encoder, and demonstrate that the model can abstract each video into much fewer visual tokens (e.g., 16) while being successful in open-ended question-answering and captioning tasks. We include a space-time attentional pooling as well as a sequential model as our temporal encoder, relying on token operations to iteratively abstract a series of frame-level tokens into a learnable memory.
There has been prior work investigating the role of pooling [16], convolutions, and cross attention layers [56, 30, 27], but study on full space-time attentional pooling or sequential model to this extent has been limited in the past. Our objective in this paper is to provide a fundamental alternative to more brute-force way of collecting all the visual tokens which have been increasing popular recently. We experimentally confirm that $16\sim 32$ video tokens abstracted by the temporal encoder is often sufficient to represent the entire video for question-answering (Figure 1).
@@ 2.1 Model architecture
We build BLIP-3-Video based on the image-based vision-language model (VLM), BLIP-3 [51]. The model architecture is composed of the following four components: (1) the vision encoder (ViT) taking each frame input, (2) the frame-level tokenizer to reduce the number of tokens, (3) the temporal encoder to build video-level token representations, and (4) the autoregressive LLM generating output text captions based on such video tokens and text prompt tokens. Figure 2 shows an overview.
@@
We use Phi-3 [1] as our LLM backbone taking such video tokens in addition to the text prompt tokens. This enables the model to take text+video as an input and generate text sentences as an output.
@@ 2.2 Temporal encoders
A temporal encoder is a function of tokens, taking $N\cdot T$ tokens as an input and returning $M$ tokens as an output: $x_{1,\ldots,M}=f(v_{(1,1),\ldots,(N,T)})$. We explore different types of encoders as part of our model. The simplest form of the temporal encoder will be temporal pooling, e.g., summating per-frame tokens over time: $x_{1,\ldots,M}=\left\{\sum_{t}(v_{(j,t)})\right\}_{j=1}^{M}$ where $M$ is restricted to be identical to $N$, which was also used in [31]. Another possible implementation would be the use of a temporal Transformer, modeling the entire token sequence and selecting the last $m$ tokens as in Mirasol3B [33]:
@@
In addition to these temporal encoders, we explore two important temporal encoders considering space-time nature of tokens: spatio-temporal attentional pooling and sequential models (Figure 3).
Spatio-temporal attentional pooling: Attentional pooling allows learnable ‘soft selection’ of multiple tokens given a larger set of tokens. It was developed for Transformers (e.g., Perceiver [13] and TokenLearner [37]), and also used in earlier foundation models (e.g., CoCa [52]) for images. In our model, we use TokenLearner [37], which explicitly serves as our space-time aware temporal encoder. Unlike previous per-frame use of pooling where spatial pooling and temporal pooling are applied separately [31], our temporal encoder directly takes all $N\cdot T$ tokens and ‘learns’ to soft-select $M$ informative tokens spatio-temporally. Here, $N$ tokens could be viewed as spatial representations of a frame and we have $T$ of them, suggesting it is a spatio-temporal representation selection.
@@
2.3 Grouped tokens sequential model
Motivated by the success of sequential models in representing series of tokens such as Mamba [12] and Token Turing Machines [38], we design a new class of sequential models for videos and take advantage of it as our temporal encoder. The sequential models are capable of taking any number of inputs to generate a fixed number of output tokens, making it suitable to build a video-level token representation (e.g., $M=32$ regardless the number of frames).
The main distinction between our new sequential models and the conventional sequential models is that our sequential model maintains a grouped memory, separately processing and maintaining different visual features. It is in a way analogous to the operational difference in the standard convolution vs. the grouped convolution, but it is done not over channels but over tokens and they are processed with multiple time steps.
Let $F$ be the main model function in a standard sequential model: $h_{i}=F(h_{i-1},v_{i})$ where $h_{i}$ is the memory token and the index $i$ ranges from 1 to $N\cdot T$ (i.e., the total number of tokens) in the entire video. In the new grouped formulation, we instead maintain the set of tokens at time step $t$ as:
$H_{t}=\bigl{\{}F(h_{(j,t-1)},v_{(j,t)})\bigr{\}}_{j=1}^{N}$ (3)
where $j$ is the visual token index within each frame. That is, we enforce the sequential model to focus on its group (specified with $j$) and maintain a ‘set’ of memory tokens for every time step $t$: $H_{t}$.
We tried both the grouped version of Mamba and TTM in our implementation as temporal encoder architectures, finding that our grouped tokens sequential model based on TTM functions better. In its instantiation, we first extend TTM by adding time-stamped positional encodings to embed the frame index of each token in the latent space. This enables the tokens in the ‘memory’ of TTM to preserve the temporal ordering information, which we found to be crucial when representing complicated or long video scenes.
Next, we implement Equation 3, while using the Read and Write operations of the original TTM. This follows our new grouped sequential model formulation, and the TTM now maintains a separate memory of size $G=4$ for each of $N=128$ tokens over time. That is, it maintains a ‘grouped’ memory, which we found to better preserve scene details. The memory is maintained to have the size of $N\cdot G$, and the final output from the sequence model is attentionally pooled from the final memory to give $M$ tokens.
In our experiments, we confirm that the proposed class of sequential models perform significantly better than the conventional sequential models like TTMs.
@@ 2.4 Training recipe
BLIP-3-Video follows a three-stage curriculum learning: (1) image caption pretraining, (2) video caption pretraining, and (3) video instruction tuning. In all its training we freeze the vision encoder, only training the model parameters after the vision encoder. First, we directly use the pretrained weights from BLIP-3 [51]. BLIP-3 is for images and it does not contain weights for the temporal encoder, so we randomly initialize those weights.
As its 2nd stage, the model is then trained on LLaVA-Hound-DPO’s video caption data [57], featuring over 900k video captions. Instead of directly using the text captions provided in LLaVA-Hound-DPO, we used GPT-4 to rephrase such text captions so that they become more GPT-style captions.
Finally, we tuned the model using a mix of video question-answering datasets, including VideoChatGPT’s 99k-sample video instruction tuning data [31], along with the training splits of the MSVD-QA [48], MSRVTT-QA [48], ActivityNet-QA [53], TGIF-QA [14], and NExT-QA [47] datasets, which contain 30k, 149k, 32k, 71k, and 34k samples, respectively. For TGIF-QA, we only used the training data associated with the Repeating Action and State Transition tasks. In our video instruction tuning recipe, we employ both open-ended and multiple-choice video QA formats for TGIF-QA and NExT-QA. For the open-ended video QA training data sourced from the MSVD-QA, MSRVTT-QA, TGIF-QA, and NExT-QA training sets, we used GPT-3.5 to rephrase the original single-word or single-phrase answer into a natural language sentence, providing the question in the LLM prompt context. For open-ended TGIF-QA and NExT-QA, we also double the sample size by using both the original short-phrase answers and the rephrased sentence-based answers. In addition, we added a filtered version of the Mira caption dataset [17] for our video instruction tuning. That is, we are using both video question-answering and video captioning for our final training. We excluded captions for Mira videos longer than one minute, totaling 935k video caption samples.
@@ 3 Experiments and Results
Model implementation details. We share the model details with BLIP-3 (4B), except that BLIP-3-Video has the new temporal encoder component in its architecture. This model takes the video with the input resolution of 384$\times$384, using SigLIP encoder to map it to 729 tokens per frame with the channel size 1152. Perceiver-Resampler is implemented with multiple cross-attention layers with the same channel dim, which is then given to the temporal encoder.
TokenLearner serving as the spatio-temporal attentional pooling was implemented using a MLP as the attention function. The size of its inner dim was the number of target tokens * 2. The grouped sequential model temporal encoder was implemented using 4 Transformer layers (with the channel dim of 1152) as the processor module while using TokenLearners for read/write modules. Memory size was set to $N*4=512$ tokens total. The resulting $16\sim 128$ tokens are mapped to the channel dim of 3072, before given to the LLM (Phi-3).
@@
VideoChat [25] 7B 32 56.3 / 2.8 45.0 / 2.5 - / 2.2 34.4 / 2.3
Video-LLaMA [56] 7B 32 51.6 / 2.5 29.6 / 1.8 12.4 / 1.1 - / -
Video-ChatGPT [31] 7B 264+ 64.9 / 3.3 49.3 / 2.8 34.2 / 2.8 51.4 / 3.0
Chat-UniVi [16] 7B 112 69.3 / 3.7 55.0 / 3.1 46.1 / 3.3 69.0 / 3.8
LLaMA-VID [27] 7B 32 69.7 / 3.7 57.7 / 3.2 47.4 / 3.3 -
LLaMA-VID [27] 13B 32 70.0 / 3.7 58.9 / 3.3 47.5 / 3.3 -
Video-LLaVA [28] 7B 2048 71.8 / 3.9 59.2 / 3.5 45.3 / 3.3 70.0 / 4.0
MiniGPT4-Video [3] 7B 2880+ 73.9 / 4.1 59.7 / 3.3 46.3 / 3.4 72.2 / 4.1
PLLaVA [49] 7B 576+ 76.6 / 4.1 62.0 / 3.5 56.3 / 3.5 77.5 / 4.1
SlowFast-LLaVA [50] 7B 3680 79.1 / 4.1 65.8 / 3.6 56.3 / 3.4 78.7 / 4.2
LLaVA-Hound-DPO [57] 7B 2048 80.7 / 4.1 70.2 / 3.7 - / - 61.4 / 3.5
LLaVA-OneVision [43] 7B 1568 72.9 / 3.9 57.8 / 3.4 55.3 / 3.6 41.1 / 3.1
Tarsier [43] 7B 4608+ 77.0 / 4.1 62.0 / 3.5 59.5 / 3.6 79.2 / 4.2
Tarsier [43] 7B 4608 74.4 / 4.0 59.1 / 3.4 54.3 / 3.5 - / -
PLLaVA [49] 34B 576+ 79.9 / 4.2 68.7 / 3.8 60.9 / 3.7 80.6 / 4.3
LLaVA-NeXT-Video [23] 32B 1152 73.6 / 4.0 56.8 / 3.4 58.4 / 3.6 73.5 / 4.1
Tarsier [43] 34B 4608+ 80.3 / 4.2 66.4 / 3.7 61.6 / 3.7 82.5 / 4.4
Tarsier [43] 34B 4608 79.3 / 4.1 62.2 / 3.5 61.5 / 3.7 - / -
@@
3.1 Public benchmarks
We conducted experiments measuring video question-answering accuracies on multiple public datasets. This includes open-ended answer generation tasks like MSVD-QA, as well as multiple choice questions like NExT-QA and MVBench [26]. We follow their standard settings in all cases.
With the temporal encoder, BLIP-3-Video was able to retain the performance with much fewer tokens, which we discuss more in the following subsection. Our results suggest that not too many visual tokens are really necessary to be successful on these open-ended question answering benchmarks, as long as we have an effective temporal encoder.
Table 2: Comparison against reported numbers of other models on NExT-QA.
Method Size #tokens Acc.
LangRepo [18] 7B 3136+ 54.6
LangRepo [18] 12B 3136+ 60.9
Tarsier [43] 7B 4608+ 71.6
LLoVi [55] 157B 1000s 67.7
IG-VLM [19] 34B 1536+ 70.9
VideoAgent [44] GPT-4 2091+ 71.3
VideoTree [45] GPT-4 3978+ 73.5
Tarsier [43] 34B 4608+ 79.2